Maleic acid is an organic compound composed of 41.39% C, 3.47% H , and the rest oxygen. If 0.168 mole of maleic acid has a mass of 19.5 g, what are the empirical and molecular fomulas of maleic acid?
1
Expert's answer
2019-10-01T03:23:30-0400

Number of moles of maleic acid"=0.168mol =\frac{19.5g}{M\frac{g}{mol}}"

So molar mass of maleic acid"=M=\frac{19.5}{0.168}=116.07\frac{g}{mol}"

Let atoms of "C" be "x"

Let atoms of "H" be "y"

Let atoms of "O" be "z"

"12\times x=0.4139\times116.07"

"1\times y=0.0347\times 116.07"

"16\times z=0.5514\times116.07"

"x\approx4"

"y\approx4"

"z\approx4"

Hence molecular formula of maleic acid is "C_4H_4O_4"

And the empirical formula is "CHO"